### v1

@@ Abstract
The ability of zero-shot translation emerges when we train a multilingual model with certain translation directions; the model can then directly translate in unseen directions. Alternatively, zero-shot translation can be accomplished by pivoting through a third language (e.g., English). In our work, we observe that both direct and pivot translations are noisy and achieve less satisfactory performance. We propose EBBS, an ensemble method with a novel bi-level beam search algorithm, where each ensemble component explores its own prediction step by step at the lower level but they are synchronized by a “soft voting” mechanism at the upper level. Results on two popular multilingual translation datasets show that EBBS consistently outperforms direct and pivot translations as well as existing ensemble techniques. Further, we can distill the ensemble’s knowledge back to the multilingual model to improve inference efficiency; profoundly, our EBBS-based distillation does not sacrifice, or even improves, the translation quality.
@@ 1 Introduction
Machine translation is a widely applicable NLP task that translates a text from a source language to a target language Brown et al. (1990); Bahdanau et al. (2015). The Transformer architecture Vaswani et al. (2017) and pretrained large language models Radford et al. (2019); Raffel et al. (2020); Lewis et al. (2020) have largely improved translation performance, especially in the supervised setting, where a model can learn from large volumes of parallel corpora. However, machine translation remains challenging for low-resource languages, because there are not enough data for large neural networks to learn these languages.
@@
We specifically focus on multilingual translation in the zero-shot setting, where the system is required to translate between unseen language pairs. Since collecting parallel data and training individual models for every translation pair are prohibitively expensive, it is common to build a single multilingual system Johnson et al. (2017); Fan et al. (2021) that can perform translation for all language pairs, most of which are zero-shot translation directions with few exceptions (e.g., English). These models work by prepending a language-indicator token, and zero-shot ability emerges as the model generalizes from trained language pairs to unseen ones (Liu et al., 2021; Wicks and Duh, 2022).
@@
The main drawback of such a multilingual model is that they are noisy in the zero-shot setting due to the lack of supervision, and as a result, they tend to generate less satisfactory outputs (Zhang et al., 2020; Liu et al., 2021). Alternatively, zero-shot translation can be performed by pivoting (Wu and Wang, 2007, 2009), where the model first translates the input into a high-resource language such as English, which is then translated to the target language. However, pivoting requires two translation steps, often leading to an accumulation of errors Babych et al. (2007); Gu et al. (2019).
In this paper, we propose a multi-pivot ensemble approach, where we design a novel ensemble method to aggregate direct translation and the predictions made by pivoting through different languages. Our ensemble approach overcomes the drawback of direct and pivot translations as it can utilize these multiple weak models, yielding high-quality output text despite lacking parallel data.
@@
Building an ensemble for text generation is nuanced as it involves a sequence of word predictions. Word-level ensembles aggregate predictions at each generation step, which is usually achieved by averaging the predicted probabilities Sennrich et al. (2016a); Freitag et al. (2017); Shanbhogue et al. (2023). This may not be ideal for zero-shot translation as the predictions are too noisy, making the averaged probabilities overly smooth. On the other hand, minimum Bayes risk decoding (Bickel and Doksum, 2015, MBR) can be considered a sequence-level voting ensemble, but it is only able to select from weak and noisy candidates given by the direct and pivot translations.
@@
To this end, we propose an ensemble decoding algorithm with bi-level beam search (EBBS). Our EBBS performs two levels of beam search at each generation step: at the lower level, beam search is applied individually to each ensemble component; at the upper level, the ensemble maintains a shared beam by voting and synchronizing the candidates (sub-sequences) in individual beams. Unlike word-level ensembles Freitag et al. (2017); Shanbhogue et al. (2023), EBBS does not average the predicted distributions, encouraging individual predictors to explore their own preferences; unlike sequence-level ensembles Kobayashi (2018); Eikema and Aziz (2020), EBBS does not select from a candidate set, and thus is more flexible since voting is performed throughout the generation process.
@@
We conducted experiments on IWSLT (Cettolo et al., 2017) and Europarl (Koehn, 2005), two popular multilingual datasets for zero-shot machine translation. Results show that EBBS can generate high-quality translations and outperform existing ensemble techniques. In addition, we used EBBS-generated outputs as the distillation signal to further finetune our multilingual translation model, which largely improves the efficiency while retaining the performance or even surpassing EBBS on Europarl.
@@ 2 Related Work
Zero-shot machine translation. With monolingual data, it is possible to accomplish zero-shot machine translation by training language-specific autoencoders that share the same latent space (Lample et al., 2018a, b; Mohiuddin and Joty, 2020). In a multilingual setting, zero-shot ability emerges when a single model is trained with multiple translation directions, allowing it to generalize to unseen directions (Dabre et al., 2020; Ranathunga et al., 2023; Johnson et al., 2017; Fan et al., 2021).
A key research direction for multilingual translation is to alleviate task interference among different translation directions (Zaremoodi et al., 2018; Wang et al., 2020). Researchers have developed various parameter-sharing schemes Baziotis et al. (2022); Chronopoulou et al. (2023) and applied optimization-based methods to locate language-specific neurons Wang and Zhang (2022); He et al. (2023); such a model is able to better learn individual translation directions while pooling the knowledge of all languages together.
In our work, we use a standard Transformer model following Johnson et al. (2017) and Liu et al. (2021). Our proposed ensemble algorithm EBBS is compatible with the above methods, as it is agnostic to model architectures.
@@
Ensemble learning. Ensemble methods integrate multiple machine learning systems to form a stronger one (Dong et al., 2020; Yang et al., 2023). Bagging, a classic ensemble technique, works by training multiple models with different portions of data and combining their predictions by averaging or voting (Breiman, 1996; Bühlmann and Yu, 2002). Another popular ensemble approach is boosting, where multiple models are trained sequentially, with each subsequent model focusing on the mistakes of the previous ones (Schapire, 2003; Hastie et al., 2009; Natekin and Knoll, 2013). Unfortunately, bagging and boosting are not compatible with our setting, because we build a self-ensemble with only a single model. Alternatively, stacking combines the outputs by training a meta-model (Wolpert, 1992; Ganaie et al., 2022), but this does not apply in our zero-shot setting either because we do not have groundtruth signal to train the meta-model.
These ensemble techniques nevertheless may be applied to supervised text generation (Freitag et al., 2017; Kobayashi, 2018; Hendy et al., 2021). That being said, they are not ideal as they do not take advantage of the sequential nature of sentences.
@@
Unlike previous work, our EBBS performs bi-level beam search, allowing each component to explore its own predictions at every step but synchronizing all individuals by a “soft voting” mechanism. Our approach is specifically suited to the sequence generation process.
@@ 3 Approach
In this section, we first explain our ensemble components in §3.1. We then propose EBBS, a novel ensemble decoding algorithm in §3.2. Finally, we describe in §3.3 knowledge distillation with EBBS-decoded outputs to improve efficiency.
@@ 3.1 Ensemble Components
Specifically, our multilingual model is an encoder–decoder Transformer with a byte pair encoding tokenizer (Sennrich et al., 2016b) shared among all languages. The encoder can capture the semantics of tokens in any language, whereas the decoder translates the encoded text into the desired language based on a target-language indicator Johnson et al. (2017); Fan et al. (2021).
We follow the standard English-centric training (Johnson et al., 2017; Liu et al., 2021), where the multilingual model is trained with parallel data with English on one of the sides (e.g., German-to-English and English-to-Romanian). The zero-shot ability emerges during such training, and the model is able to translate directly between unseen language pairs (e.g., German-to-Romanian), known as direct translation Dabre et al. (2020); Ranathunga et al. (2023). An alternative approach is pivot translation, where the multilingual model performs two translations using a high-resource language as a pivot (e.g., first translating German to English, and then English to Romanian).
However, both direct and pivot translations have major weaknesses: the quality of direct translation tends to be low due to the lack of parallel data, whereas pivot translation suffers from an error accumulation as it requires two translation steps (Babych et al., 2007; Gu et al., 2019).
In this paper, we would like to build a self-ensemble of the multilingual model to boost translation quality, where direct and pivot translations are our ensemble components.
However, standard ensemble methods such as averaging and voting may not work well for text generation. Voting, for example, chooses the most voted prediction, but in text generation, the components’ votes often do not share anything in common, because there could be tens of thousands of tokens in the vocabulary. An averaging ensemble, on the other hand, averages the predicted distributions of all components, potentially leading to an overly smooth distribution. As reported in a pilot study in the appendix of Fan et al. (2021), a multilingual averaging ensemble only achieves a small improvement over direct translation. Overall, ensemble methods have not been widely explored for zero-shot multilingual translation in previous literature.
@@ 3.2 Our Proposed EBBS Algorithm
Figure 1: Overview of our EBBS algorithm.
@@ 1 $H\leftarrow\emptyset$triangleright candidate outputs
2 $B_{0}\leftarrow\{\langle\textsc{bos},1\rangle\}$triangleright upper-level beam
@@
4       $\triangleright$ lower level: individual beam search
5       for $\langle\mathbf{y}_{1:t-1},p\rangle\in B_{t-1}$ do
@@ 4       $\triangleright$ lower level: individual beam search
7                   $B^{\prime}_{t,k}\leftarrow\emptyset$triangleright lower-level beam
@@
10                         $B^{\prime}_{k}.\operatorname{add}(\langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},p\cdot p^{\prime}\rangle)$
11                  $B^{\prime}_{t,k}\leftarrow B^{\prime}_{t,k}.\operatorname{top}(Z)$
@@
13      $\triangleright$ upper level: beam synchronization
@@ 14       $D\leftarrow\text{empty dictionary}$
16             for $\langle\mathbf{y},p\rangle\in B^{\prime}_{t,k}.\operatorname{top}(Z)$ do
@@ 24       $\triangleright$ check for termination
25       for $\langle\mathbf{y},p\rangle\in B_{t}$ do
@@
To this end, we propose an ensemble with bi-level beam search (EBBS), a novel decoding framework that enables different ensemble components to collaborate and vote on each other’s partial generations with two levels of beam search.
For the 00th decoding step, EBBS initializes a shared, upper-level beam $B_{0}=\langle\textsc{bos},1\rangle$, suggesting that a sequence is forced to start with a special token bos with probability $1$.
For step $t$, each ensemble component performs a lower-level beam search individually, based on the prefixes in the last step’s shared beam $B_{t-1}$:
$B_{t,k}^{\prime}=$ $\operatorname{top-}\!Z\{\ \langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},\ \ p\cdot p_{k}(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})\rangle:$
$\hskip 37.5807pt\langle\mathbf{y}_{1:t-1},p\rangle\in B_{t-1},\ \ \mathrm{y}\in V\}$ (1)
At the upper level, EBBS synchronizes the individual beams $B_{t,k}^{\prime}$, for $k=1,\cdots,K$, into a single, shared, upper-level beam through a soft voting mechanism, where the candidate set $C_{t}$ is the union of the sequences in individual beams:
$C_{t}=\bigcup\nolimits_{k}\{\mathbf{y}:\langle\mathbf{y},p\rangle\in B_{t,k}^{\prime}\}$ (2)
Then, we evaluate each candidate in $C_{t}$ and compute its overall vote as the sum of the probabilities.
[…]
In this way, the upper level synchronizes every ensemble component with the shared beam $B_{t}$ for the next step of generation.
Intuitively, our voting scheme gives an ensemble component $Z$-many votes, each weighted by the predicted probability. The votes (probabilities) are then tallied (summed) for each candidate to form the upper-level beam. Our bi-level beam search terminates when we have $Z$-many terminated sequences in the shared beam, and returns the sequence with the highest score [^1] For selecting the final output, we follow standard implementations and normalize the joint probabilities by length, i.e., taking the geometric mean of step-wise probabilities (Wolf et al., 2019; Ott et al., 2019). Otherwise, beam search algorithms are often biased towards short sequences. as the ensemble output. We provide the detailed pseudocode for EBBS in Algorithm 1 and an illustration in Figure 1.
Discussion. As shown in Algorithm 2 (Appendix A), traditional beam search keeps a fixed-size beam of high-likelihood partial sequences. It is tempting to combine such vanilla beam search with multiple predictors by directly averaging their probabilities $p(\mathbf{y}|\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}p_{k}(\mathbf{y}|\mathbf{x})$ as the score for beam search (Line 8, Algorithm 2), which has been experimented in previous work Sennrich et al. (2016a); Shanbhogue et al. (2023).
However, our intuition suggests that such an approach may suffer from the over-smoothing problem Wei et al. (2019); Wen et al. (2023b): when multiple translations (known as modes) are plausible given an input, the ensemble process will overly smooth out the modes by probability averaging.
By contrast, EBBS allows each ensemble component to explore its own mode (Lines 4–11, Algorithm 1). In Figure 1, for example, the top sequence yields two plausible next tokens, suggested by each component in the lower level. Their probabilities are not smoothed out in our approach, unlike averaging ensembles. The upper level performs soft voting (Lines 12–19, Algorithm 1) so as to maintain tractable inference.
3.3 EBBS-Based Distillation
To improve inference efficiency, we may perform knowledge distillation based on the outputs of our EBBS algorithm. In particular, we follow Kim and Rush (2016) and apply a sequence-level knowledge distillation loss, treating the output $\hat{\mathbf{y}}$ of our ensemble (called a teacher) as the pseudo-groundtruth for finetuning the multilingual translation model (called a student):
@@ 3.3 EBBS-Based Distillation
Our self-distillation is an ensemble-then-distill process. This differs from a straightforward practice of multi-teacher distillation, where the student learns from the union of teachers’ outputs Wu et al. (2021). The commonly applied cross-entropy loss is known to yield overly smooth distributions Wen et al. (2023a, b), and the problem becomes more severe with multiple teachers, leading to unsatisfactory performance of union distillation Shayegh et al. (2024). By contrast, our approach provides the student with a consolidated pseudo-groundtruth translation for an input, causing less confusion during the distillation process especially when teachers disagree with each other.

### v2

@@ Abstract
The ability of zero-shot translation emerges when we train a multilingual model with certain translation directions; the model can then directly translate in unseen directions. Alternatively, zero-shot translation can be accomplished by pivoting through a third language (e.g., English). In our work, we observe that both direct and pivot translations are noisy and achieve less satisfactory performance. We propose EBBS, an ensemble method with a novel bi-level beam search algorithm, where each ensemble component explores its own prediction step by step at the lower level but all components are synchronized by a “soft voting” mechanism at the upper level. Results on two popular multilingual translation datasets show that EBBS consistently outperforms direct and pivot translations, as well as existing ensemble techniques. Further, we can distill the ensemble’s knowledge back to the multilingual model to improve inference efficiency; profoundly, our EBBS-distilled model can even outperform EBBS as it learns from the ensemble knowledge.
GitHubhttps://github.com/MANGA-UOFA/EBBS
@@ 1 Introduction
Machine translation is a widely applicable NLP task that aims to translate a text from a source language to a target language (Brown et al. 1990; Bahdanau, Cho, and Bengio 2015). The Transformer architecture (Vaswani et al. 2017) and pretrained large language models (Radford et al. 2019; Lewis et al. 2020) have largely improved translation performance, especially in the supervised setting (Raffel et al. 2020), where a model can learn from large volumes of parallel corpora. However, machine translation remains challenging for low-resource languages, because there are not enough data for large neural networks to learn these languages (Radford et al. 2019; Muennighoff et al. 2023).
@@
We specifically focus on multilingual translation in the zero-shot setting, where the system is required to translate between unseen language pairs. Since collecting parallel data and training individual models for every translation pair are prohibitively expensive, it is common to build a single multilingual system (Johnson et al. 2017; Fan et al. 2021) that can perform translation for all language pairs, most of which are zero-shot translation directions that do not involve a high-resource language (e.g., English). These models work by prepending a language-indicator token; the zero-shot translation ability emerges as the model generalizes from trained language pairs and is able to perform direct translation for unseen ones (Liu et al. 2021; Wicks and Duh 2022). The main drawback of such multilingual models is that they are noisy in the zero-shot setting due to the lack of supervision, and as a result, they tend to generate low-quality translations (Zhang et al. 2020; Liu et al. 2021).
@@
Alternatively, zero-shot translation can be performed by pivoting (Wu and Wang 2007, 2009), where the model first translates the input into a high-resource language such as English, which is then translated to the target language. However, pivoting requires two translation steps, often leading to an accumulation of errors (Babych, Hartley, and Sharoff 2007; Gu et al. 2019).
@@
In this paper, we propose an ensemble approach that aggregates direct and pivot translations in order to build a stronger multilingual translation model from weak ones. Building an ensemble for text generation is nuanced as it involves a sequence of word predictions. Word-level ensembles aggregate predictions at each generation step, which is usually achieved by averaging the predicted probabilities (Sennrich, Haddow, and Birch 2016a; Freitag, Al-Onaizan, and Sankaran 2017; Shanbhogue et al. 2023). This may not be ideal for zero-shot translation as the predictions are too noisy, making the averaged probabilities overly smooth. On the other hand, minimum Bayes risk decoding (MBR) (Bickel and Doksum 2015) can be considered a sequence-level voting ensemble, but existing MBR methods are only able to select from weak and noisy candidates given by the direct and pivot translations.
@@
To this end, we propose an ensemble decoding algorithm with bi-level beam search (EBBS). Our EBBS performs two levels of beam search at each generation step: at the lower level, beam search is applied individually to each ensemble component; at the upper level, the ensemble maintains a shared beam by voting and synchronizing the candidates (sub-sequences) in lower-level beams. Unlike word-level ensembles (Freitag, Al-Onaizan, and Sankaran 2017; Shanbhogue et al. 2023), EBBS does not average the predicted distributions, encouraging individual predictors to explore their own preferences; unlike sequence-level MBR ensembles (Kobayashi 2018; Eikema and Aziz 2020), EBBS does not select from a candidate set, and thus is more flexible since votings are performed throughout the generation process.
@@
We conducted experiments on IWSLT (Cettolo et al. 2017) and Europarl (Koehn 2005), two popular multilingual datasets for zero-shot machine translation. Results show that EBBS can generate high-quality translations and outperform existing ensemble techniques. In addition, we used EBBS-generated data for distillation to further improve the multilingual model. The experiment shows that such a distilling process encourages the model to learn from high-quality translations produced by EBBS, allowing it to outperform EBBS with no inference overhead compared with direct translation.
@@ 2 Related Work
Machine translation. In NLP, machine translation is a long-standing task that aims to rewrite text from one language to another without changing the meaning. Traditional research in translation has been mainly centered on the supervised setting, utilizing manually crafted rules (Forcada et al. 2011; Dugast, Senellart, and Koehn 2007) and statistical methods (Brown et al. 1990; Koehn 2009); more recently, neural machine translation systems have considerably improved the performance (Vaswani et al. 2017; Raffel et al. 2020). However, translation remains challenging for low-resource languages, where neural models do not have enough parallel data to train on.
Translation for low-resource languages largely relies on zero-shot techniques, where no parallel text is available for a particular translation direction. In general, zero-shot translation can be accomplished in a monolingual or multilingual setting. With monolingual data, the most common approach is to build language-specific autoencoders that share the same latent space of semantics; translation is then achieved by plugging in the decoder of the desired language (Lample et al. 2018a, b; Mohiuddin and Joty 2020).
In this paper, we focus on the multilingual setting, where one model can translate between multiple languages (Dabre, Chu, and Kunchukuttan 2020). Usually, parallel texts only exist for a high-resource language such as English, leaving translations between low-resource languages zero-shot (e.g., Italian to Dutch) (Johnson et al. 2017; Fan et al. 2021). In this setting, the most common approach is to train the multilingual model on English-centric data, and the zero-shot translation ability naturally emerges during the training process (Johnson et al. 2017; Scao et al. 2022).
A key challenge for multilingual models is task interference, where too many languages tend to degrade model performance (Zaremoodi, Buntine, and Haffari 2018; Wang, Lipton, and Tsvetkov 2020). As a result, research in this direction has been alleviating such interference by developing various parameter-separation schemes (Baziotis et al. 2022; Chronopoulou, Stojanovski, and Fraser 2023) and using gradient-based methods to update language-specific parameters (Wang and Zhang 2022; He et al. 2023). In our work, we use a standard Transformer model following Johnson et al. (2017) and Liu et al. (2021). Our proposed ensemble algorithm EBBS is compatible with the above approaches, as it is agnostic to model architectures.
@@
Ensemble methods. In a model ensemble, multiple machine learning systems are integrated so as to form a stronger one (Dong et al. 2020; Yang, Lv, and Chen 2023). Bagging, a classic ensemble technique, works by training multiple models with different portions of data and combining their predictions through averaging or voting (Breiman 1996; Bühlmann and Yu 2002). Another popular ensemble approach is boosting, where different models are trained sequentially, with each subsequent model addressing the mistakes of the previous ones (Schapire 2003; Hastie et al. 2009; Natekin and Knoll 2013). Unfortunately, bagging and boosting are not compatible with our setting, because we build an ensemble with a single model. Alternatively, stacking combines the outputs by training a meta-model (Wolpert 1992; Ganaie et al. 2022), but this does not apply to our zero-shot setting either because we do not have groundtruth signals to train the meta-model. Even though these ensemble techniques may be directly applied to supervised generation (Freitag, Al-Onaizan, and Sankaran 2017; Kobayashi 2018; Hendy et al. 2021), they are not ideal as they do not take advantage of structured prediction. Our recent work has addressed the ensemble of tree structures (Shayegh et al. 2024; Shayegh, Wen, and Mou 2024; Shayegh et al. 2025), and in this paper we focus on text generation.
@@
Unlike previous work, our EBBS performs bi-level beam search, exploring different components’ own predictions and synchronizing them by a “soft voting” mechanism at every step. Our approach is specifically suited to the sequence generation process.
@@ 3 Approach
In this section, we first explain our ensemble components in §3.1. In §3.2, we propose EBBS, a novel ensemble decoding algorithm. Finally, we describe in §3.3 knowledge distillation with EBBS-decoded outputs for efficiency considerations.
@@ 3.1 Ensemble Components
Specifically, our multilingual model is an encoder–decoder Transformer with a byte pair encoding tokenizer (Sennrich, Haddow, and Birch 2016b) shared among all languages. The encoder can capture the semantics of tokens in different languages, whereas the decoder translates the encoded text into the desired language based on a target-language indicator token (Johnson et al. 2017; Fan et al. 2021).
We follow the standard English-centric training (Johnson et al. 2017; Liu et al. 2021), where the multilingual model is trained using parallel data with English on one side (e.g., German-to-English and English-to-Romanian). As mentioned in §1, the zero-shot ability emerges during such training, and the model is able to perform direct translation between unseen language pairs (e.g., German-to-Romanian) (Dabre, Chu, and Kunchukuttan 2020; Ranathunga et al. 2023). An alternative approach is pivot translation, where the multilingual model performs two translations using a high-resource language as a pivot (e.g., first translating German to English, and then English to Romanian).
However, both direct and pivot translations have major weaknesses: the quality of direct translation tends to be low due to the lack of parallel data, whereas pivot translation suffers from error accumulation as it requires two translation steps (Babych, Hartley, and Sharoff 2007; Gu et al. 2019).
In this paper, we would like to build an ensemble of direct and pivot translations to boost translation quality, where each translation path results in an ensemble component. Commonly used ensemble methods such as averaging and voting may not work well for text generation. Voting, for example, chooses the most voted prediction, but in text generation, the components’ votes often do not share anything in common, because there could be tens of thousands of tokens in the vocabulary. An averaging ensemble, on the other hand, averages the predicted distributions of all components, potentially leading to an overly smooth distribution. Despite early success by Razmara and Sarkar (2013), more recent studies report marginal or negative improvement for multi-pivot averaging ensemble (Fan et al. 2021; Gaikwad et al. 2024; Mohammadshahi, Vamvas, and Sennrich 2024).
@@ 3.2 Our Proposed EBBS Algorithm
Figure 1: Illustration of our EBBS algorithm.
We propose an ensemble with bi-level beam search (EBBS), a novel decoding algorithm that enables different ensemble components to collaborate and vote on each other’s partial generations with two levels of beam search.
For the 00th decoding step, EBBS initializes the upper-level beam by $\overline{B}_{0}=\langle\textsc{bos},1\rangle$, suggesting that a sequence is forced to start with a special beginning-of-sequence token bos with probability $1$.
For step $t$, each ensemble component performs lower-level beam search individually, based on the prefixes in the last step’s shared beam $\overline{B}_{t-1}$:
[…]
$\langle\mathbf{y}_{1:t-1},p\rangle\in\overline{B}_{t-1},\ \ \mathrm{y}\in V\ \}$ (1)
At the upper level, EBBS synchronizes the lower-level individual beams $\underline{B}_{t,k}$, for $k=1,\cdots,K$, into a shared, upper-level beam through a soft-voting mechanism, where the candidate set $C_{t}$ is the union of the sequences in lower-level beams:
$C_{t}=\bigcup\nolimits_{k}\{\mathbf{y}:\langle\mathbf{y},p\rangle\in\underline{B}_{t,k}\}$ (2)
We evaluate each candidate in $C_{t}$ and compute its overall vote as the sum of the probabilities.
[…]
In this way, the upper level synchronizes all ensemble components with the shared beam $\overline{B}_{t}$ for the next step of generation.
Intuitively, our voting scheme gives an ensemble component $Z$-many votes, each weighted by the predicted probability. The votes (probabilities) are then tallied (summed) for each candidate to form the upper-level beam. Our bi-level beam search terminates when we have $Z$-many terminated sequences in the shared beam, and returns the sequence with the highest score [^1] For selecting the final output, we follow standard implementations and normalize the joint probability by length, i.e., taking the geometric mean of step-wise probabilities (Wolf et al. 2019; Ott et al. 2019). Otherwise, beam search algorithms are often biased towards short sequences (Meister, Cotterell, and Vieira 2020). as the ensemble output. We provide the detailed pseudocode for EBBS in Algorithm 1 and an illustration in Figure 1.
Discussion. Traditional beam search keeps a fixed-size beam of high-likelihood partial sequences. To build an ensemble with multiple predictors, it is tempting to directly average their probabilities $p(\mathbf{y}|\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}p_{k}(\mathbf{y}|\mathbf{x})$ as the score for beam search, which has been experimented in previous work (Sennrich, Haddow, and Birch 2016a; Shanbhogue et al. 2023).
However, our intuition suggests that such an approach may suffer from the over-smoothing problem (Wei et al. 2019; Wen et al. 2023b): when multiple translations (known as modes) are plausible given an input, the ensemble process will overly smooth out the modes by probability averaging.
By contrast, EBBS allows each ensemble component to explore its own mode (Lines 4–11, Algorithm 1). In Figure 1, for example, the top sequence yields two plausible next tokens, suggested by each component in the lower level; their probabilities are not smoothed out in our approach, unlike averaging ensembles. The upper level performs soft voting (Lines 12–19, Algorithm 1) so as to maintain tractable inference.
3.3 EBBS-Based Distillation
@@ 1 $H\leftarrow\emptyset$triangleright candidate outputs
2 $\overline{B}_{0}\leftarrow\{\langle\textsc{bos},1\rangle\}$triangleright upper-level beam
@@
4       $\triangleright$ lower: individual beam search
5       for $\langle\mathbf{y}_{1:t-1},p\rangle\in\overline{B}_{t-1}$ do
@@ 4       $\triangleright$ lower: individual beam search
7                   $\underline{B}_{t,k}\leftarrow\emptyset$triangleright lower-level beam
@@
10                         $\underline{B}_{t,k}.\operatorname{add}(\langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},p\cdot p^{\prime}\rangle)$
11                  $\underline{B}_{t,k}\leftarrow\underline{B}_{t,k}.\operatorname{top}(Z)$
@@
13      $\triangleright$ upper: beam synchronization
@@ 14       $D\leftarrow\text{empty dictionary}$
16             for $\langle\mathbf{y},p\rangle\in\underline{B}_{t,k}$ do
@@ 24       $\triangleright$ check for termination
25       for $\langle\mathbf{y},p\rangle\in\underline{B}_{t}$ do
@@
To improve inference efficiency, we perform knowledge distillation based on the outputs of our EBBS algorithm. In particular, we follow (Kim and Rush 2016) and apply a sequence-level knowledge distillation loss, treating the output $\hat{\mathbf{y}}$ of our ensemble (serving as a teacher) as the pseudo-groundtruth for finetuning the multilingual translation model (serving as a student):
@@
Our distilling method is an ensemble-then-distill process. This differs from a straightforward practice of multi-teacher distillation, where the student learns from the union of teachers’ outputs (Wu, Wu, and Huang 2021). The commonly applied cross-entropy loss is known to yield overly smooth distributions (Wen et al. 2023a, b), and the problem becomes more severe with multiple teachers, leading to less satisfactory performance of union distillation (Shayegh et al. 2024). On the contrary, our approach provides the student with a consolidated pseudo-groundtruth translation, causing less confusion during the distillation process especially when teachers disagree.